mathslover (mathslover):

Evaluate : \(\cos 12 \cos 24 \cos 36 \cos 48 \cos 72 \cos 84\)

mathslover (mathslover):

@Mertsj

mathslover (mathslover):

I was thinking to simplify \(\cos 72\) more as the pattern was going +12 ... but 48 + 12 = 60 that broke the pattern.

mathslover (mathslover):

\(\cos 12 \cos (2(12)) \cos (3(12)) \cos (4(12)) \cos (6(12)) \cos (7(12))\)

mathslover (mathslover):

We have multiple angles.

mathslover (mathslover):

\( (\cos 12 \cos 24 \cos 48 )( \cos 36 \cos 72 \cos (180-96)\) Oh I have something useful here.

OpenStudy (anonymous):

cosx .cos(60-x) . cos(60+x) = cos3x/4

OpenStudy (anonymous):

use this here

OpenStudy (anonymous):

how about noticing 12+48=60, 12-48=-36 24+36=60, 24-36=-12

mathslover (mathslover):

\( - (\cos 12 \cos 48 \cos 24 \cos 96) (\cos 36 \cos 72) \) Hmn .. @electrokid but what would I get from there?

OpenStudy (anonymous):

my bad, its 72,

OpenStudy (anonymous):

\[\cos(12°) \cos(24°) \cos(36°) \cos(48°) \cos(72°) \cos(84°)\] \[1/64\]

OpenStudy (anonymous):

Maybe double angle for \(\cos\) would be good.

OpenStudy (anonymous):

\[ \cos12\cos48={1\over2}[\cos(12-48)-\cos(12+48)]={1\over2}\left(\cos36-{1\over2}\right)={1\over2}\cos36-{1\over4} \]

OpenStudy (anonymous):

hmmmm...

OpenStudy (anonymous):

\[ \cos24\cos36={1\over2}[\cos(24-36)-\cos(24+36)]={1\over2}\left(\cos12-{1\over2}\right)={1\over2}\cos12-{1\over4} \]

OpenStudy (anonymous):

\[ \cos(2x) = \cos^2(x)-\sin^2(x) \]

OpenStudy (anonymous):

@electrokid, it'd be nice if there were more factors of \(\cos(5(12))\) to work with... Anything we can do with the remaining \(\cos(6(12))\) and \(\cos(7(12))\)?

OpenStudy (anonymous):

I hope the product expansion of these parantheses would provide us with something \[\cos (x)\cos (2x)\cos(3x)\cos(4x)\cos(6x)\cos(7x)\]

OpenStudy (anonymous):

let θ=mean of the angles?

OpenStudy (anonymous):

\[\huge\frac{ 1 }{ 4 }sin60*\cos12*\cos24\]

OpenStudy (anonymous):

@mathslover what is the answer??

mathslover (mathslover):

I am getting it as \(\cfrac{1}{64}\)

OpenStudy (anonymous):

hmmm i got that too xD

OpenStudy (anonymous):

is there cos 60 also in the question??

mathslover (mathslover):

no!

OpenStudy (anonymous):

then how come u got 1/64 ??

OpenStudy (anonymous):

oh yes yes

OpenStudy (anonymous):

the answer is \[\huge\frac{ 1 }{ 64 }\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 64 } \approx 0.0156250\]

OpenStudy (anonymous):

the answer is 1/64

OpenStudy (anonymous):

i took cos 60 also so i got my previous answer!!

OpenStudy (anonymous):

cos(12^(o))cos(24^(o))cos(36^(o))cos(48^(o))cos(72^(o))cos(84^(o)) Take the cosine of 12^(o) to get 0.9782. (0.9782)(cos(24^(o))cos(36^(o))cos(48^(o))cos(72^(o))cos(84^(o))) Take the cosine of 24^(o) to get 0.9136. (0.9782)(0.9136)(cos(36^(o))cos(48^(o))cos(72^(o))cos(84^(o))) Take the cosine of 36^(o) to get 0.809. (0.9782)(0.9136)(0.809)(cos(48^(o))cos(72^(o))cos(84^(o))) Take the cosine of 48^(o) to get 0.6691. (0.9782)(0.9136)(0.809)(0.6691)(cos(72^(o))cos(84^(o))) Take the cosine of 72^(o) to get 0.309. (0.9782)(0.9136)(0.809)(0.6691)(0.309)(cos(84^(o))) Take the cosine of 84^(o) to get 0.1045. (0.9782)(0.9136)(0.809)(0.6691)(0.309)(0.1045) Multiply 0.9782 by 0.9136 to get 0.8936. (0.8936)(0.809)(0.6691)(0.309)(0.1045) Multiply 0.8936 by 0.809 to get 0.7229. (0.7229)(0.6691)(0.309)(0.1045) Multiply 0.7229 by 0.6691 to get 0.4837. (0.4837)(0.309)(0.1045) Multiply 0.4837 by 0.309 to get 0.1495. (0.1495)(0.1045) Multiply 0.1495 by 0.1045 to get 0.0156. (0.0156) Remove the parentheses around the expression 0.0156. 0.0156 So this is how I got it which i got 0.0156 and 1/64 is ~ 0.0156 DO YOU UNDERSTAND HOW TO DO IT NOW, EVERYONE?

mathslover (mathslover):

@some_someone took from .. software? lol

OpenStudy (anonymous):

yeah i checked it using wolfram?

mathslover (mathslover):

:)

OpenStudy (anonymous):

Yeah! Its \(\cfrac{1}{64}\) . You're right @mathslover